The length of one of the equal legs of an isosceles triangle is 8cm less than 4 times the lenghth of the base. If the perimeter is 29cm, find the lenghth of the equal legs.
Answer by stanbon(48558)

The length of one of the equal legs of an isosceles triangle is 8cm less than 4 times the lenghth of the base. If the perimeter is 29cm, find the lenghth of the equal legs.
----------------
Let the length of the base be "x" cm.
Leg = 4x-8
Other leg = 4x-8
-------------
EQUATION:
x + 2(4x-8) = 29
9x - 16 = 29
9x = 45
x = 5 cm (length of the base)
4x-8 = 12 cm (length of one of the equal legs)
==============
Cheers,
Stan H.